# thanks: Equal contributionsthanks: Equal contributions

### v2

@@ Abstract
Furthermore, some features in Fig. 3 require further explanation. First, the disappearance of half-integer Shapiro steps at zero field (Fig. 3a, c) may be attributed to the relatively small contribution of the second harmonic in the CPR and the low microwave frequency 58. In device S1, a higher microwave frequency is utilized at zero field, and a small signal of half-integer steps appears in supplementary Fig. S6d-f. No higher frequency is applied to detect half-integer steps at 0 T, considering the limitations of our microwave generator. Second, the integer Shapiro steps at 5.5 mT (3/2 $\phi_{0}$) appear weaker in Fig. 3b (or supplementary Fig. S6c) compared to those at 0 T in supplementary Fig. S6b. This is due to the application of a magnetic field at half the flux quantum, causing a reduction in the sharpness of the integer steps because of a decrease in the first harmonic of the CPR. Subsequently, the second harmonic or fractional Shapiro steps become visible at half the flux quantum (Fig. 3a), in agreement with theoretical predictions 9; 35 and experimental results 53; 24. Third, in device S2 (Fig. 3c), half-integer steps vanish at ±$\pm\phi_{0}/2$ due to the superposition of multiple SQUID patterns, which is slightly different from the single SQUID pattern observed in device S1. In the case of multiple SQUID patterns, the position of destructive interference for the first harmonic differs from that of a single SQUID pattern, leading to a notable contribution from the first harmonic and the absence of half-integer steps at ±$\pm\phi_{0}/2$ (see supplementary section VI).
Figure 4: Magnetochiral anisotropy of Ta2Pd3Te5 device S2. $\mathbf{a}$, $R_{w}$ versus $B_{z}$ at different temperatures. $\mathbf{b}$, Antisymmetric $R_{2w}$ with respect to $B_{z}$ at various temperatures. $\mathbf{c,d}$, $R_{w}$ and $R_{2w}$ curves at 20 mK. Pairs of antisymmetric peaks ($R_{2w}$) are marked by dashed lines with the same colors. $R_{2w}$ shows the similar behavior to $\Delta I_{c}$ in $\mathbf{d}$. $\mathbf{e}$, MCA coefficient $\gamma$ as a function of $B_{z}$ from 20 mK to 0.45 K. $\gamma$ decreases with increasing $B_{z}$. $\mathbf{f}$, Temperature-dependent $\gamma$. The average $\gamma$ is calculated using pairs of antisymmetric peaks in $\mathbf{d}$. $\mathbf{g}$, Summary of $\gamma^{\prime}$ as a function of $R_{2w}/R_{w}$. Ta2Pd3Te5 interferometer exhibits the huge $\gamma$ at large $R_{2w}/R_{w}$. The open (half-fulled, fulled) scatters represent the JJs (normal conductors, intrinsic superconductors).
@@
The JDE is also studied as a function of microwave power. In Fig. 3f, switching current $I_{c+}^{0th}$ and $I_{c-}^{0th}$ of the 0th Shapiro step are extracted from the Fig. 3e. The quantity $\Delta I_{c}^{0th}=I_{c+}^{0th}-\left|I_{c-}^{0th}\right|$ (orange curve, see the detail in supplementary Fig. S8) displays little variation and gradually decreases with increasing microwave power. This stable observation of the JDE under microwave irradiation represents one of the initial examples of such behavior and is a crucial feature in quantum information applications, where microwaves are commonly utilized 2.
Interfering MCA
@@
As another type of non-reciprocal behavior, MCA is measured to characterize non-reciprocal charge transport 59; 13; 25, with its strength often defined by the coefficient $\gamma=\left|\frac{2R_{2w}}{R_{w}BI}\right|$. MCA typically exhibits antisymmetric behavior in $R_{2w}$ with respect to magnetic fields. In normal conductors, MCA is a small effect ($\gamma$ $<$ 1 T-1A-1), generally due to the small SOC. However, significant enhancement of MCA is observed in noncentrosymmetric superconductors, possibly due to a large difference in the energy scale between the Fermi energy and the superconducting gap 59.
Therefore, it is quite interesting to measure MCA in a supercurrent interferometer. In Fig. 4a, b, the first- ($R_{w}$) and second-harmonic resistances ($R_{2w}$) of device S2 are measured using lock-in techniques. Unlike the linear current dependence of $\left|R_{2w}\right|$ in other systems at small current 60; 61; 62, current-dependent $R_{2w}$ behaviors in our JJs seem a little complex at different fields because of supercurrent interference (supplementary section VII). As a result, an ac current $I_{ac}$ = 500 nA is applied to measure temperature-dependent $R_{2w}$ in Fig. 4 to obtain a large $R_{2w}$ and observe relatively intact features of antisymmetric $R_{2w}$ $B_{z}$ curves. In fact, the oscillating antisymmetric behavior of $R_{2w}$ $B_{z}$ curves in Fig. 4b is different from previous reports 59; 13; 25; 54, where only a pair of antisymmetric peaks exist with respect to the magnetic field. These antisymmetric behaviors can usually provide additional information compared to SDE, including disparities between the Fermi energy and superconducting gap, the Rashba parameter and spin-orbit splitting, among other factors 59; 63. The oscillating $R_{2w}$ (red line in Fig. 4d) is analogous to the $\Delta I_{c}$ of JDE (orange line in Fig. 4d, and similar behavior of device S1 in supplementary Fig. S10-12), indicating a common underlying origin. To our best knowledge, such interfering MCA has not been observed before, and apparently reduces magnetic fields at which MCA appears to a large degree.
In Fig. 4e, an extremely large $\gamma$ = 2.2 ×$\times 10^{9}$ T-1A-1 is obtained at tiny $B_{z}$, suggesting that this is the largest value reported to date 54; 61; 62; 13; 25; 60; 64; 65. With increasing $B_{z}$ and temperature, $\gamma$ gradually decreases, as shown in Fig. 4e, f, respectively. In Fig. 4f, the average $\gamma$ is calculated using pairs of antisymmetric peaks (black and red dashed lines in Fig. 4c, d), and the largest $\gamma_{1}$ and $\gamma_{2}$ are $1.2\times 10^{9}$ and $9.6\times 10^{8}$ T-1A-1, respectively. To compare with the MCA in several typical systems, the relationship between $\gamma^{\prime}$ and $R_{2w}/R_{w}$ is summarized in Fig. 4g, where $\gamma^{\prime}$ ($=\gamma A$, where $A$ is the cross-sectional area of the JJ) is also calculated to characterize the MCA, because $\gamma$ seems to show a positive relation with $R_{2w}/R_{w}$ (supplementary Fig. S9). Although no evident relationship between $\gamma^{\prime}$ and $R_{2w}/R_{w}$ exists in different SDE/JDE systems (Fig. 4g), Ta2Pd3Te5 JJs possess not only the large $R_{2w}/R_{w}$ but also huge $\gamma^{\prime}$. MCA of the device S1 is shown in supplementary Fig. S10-12. The maximum value of $\gamma^{\prime}$ can reach $8.44\times 10^{-5}$ ($1.3\times 10^{-4}$) m2T-1A-1 for device S2 (device S1) with $A$ = $7.03\times 10^{-14}$ ($2.08\times 10^{-14}$) m2, which seems to be the largest value reported so far 64; 65. Furthermore, MCA in interferometers can be greatly enhanced theoretically by increasing the enclosed area of the SQUID loop. Here, the enhancement of $\gamma^{\prime}$ is apparently correlated to the interference of asymmetric edge supercurrents, probably due to the much larger energy difference between the Fermi energy and the mini-gap in Andreev bound states, instead of the bulk superconducting gap 59. In this scenario, the mini-gap is modulated by the flux quantum pinned in the SQUID, which is why $R_{2w}$ oscillates in conjunction with the supercurrent interference. Further investigation of MCA in such types of interference systems is warranted.
Conclusions
@@
The Ta2Pd3Te5 edge interferometer reveals a significant JDE under small out-of-plane magnetic fields, as well as extremely large MCA. Moreover, the presence of asymmetric Josephson critical current alongside fractional Shapiro steps demonstrates the importance of the higher harmonic content in the JDE. It is noted that power consumption requires attention if SDE/JDE is to be applied at low temperatures in the future, and further optimization is necessary not only regarding mechanism, but also in terms of device size. The Josephson diode efficiency can be further enhanced by concatenating interferometer loops. These findings offer a promising and novel approach to exploring JDE with significant diode efficiency, ultra-low switching power at small magnetic fields and stable JDE under microwave irradiation, rendering them potentially valuable for practical applications.
Motheds
@@
Device fabrication. Single crystals Ta2Pd3Te5 were prepared by the self-flux method 36. Ta2Pd3Te5 thin films were obtained through mechanical exfoliating bulk samples onto Si substrates with the 280-nm-thick SiO2 on top, and then coated with PMMA in a glove box at 4000 rpm for 60 s, followed by annealing at 100 C for 120 s. Multi-terminal electrical contacts were patterned after electron-beam exposure and subsequent development. Ti/Al (5 nm/60 nm) electrodes were deposited after Ar etching in order to remove the oxidized layer. After the lift-off step, the device was coated with boron nitride to prevent oxidation. The entire fabrication process took place in a nitrogen atmosphere glove box. The rectangular film, with a long side along the $b$-axis of the crystal, was easily obtained due to its quasi-1D nature 38. The dimensions of device S2 (device S1) are as follows: thickness 37 nm (16 nm), width 1.9 $\mu$m (1.3 $\mu$m), and length 245 nm (400 nm), respectively.
@@
Transport measurements under microwave radiation. The electrical transport measurements under microwave radiation were carried out using PSG-A Series Singal generator (Agilent E8254A). The main microwave frequency between 4.64 and 5.02 GHz was applied due to the large absorption effect of microwave for devices S1 and S2. The dissipation of microwave circuit existed during the measurement of device S2, while another microwave circuit for device S1 was good. So the applied power of microwave was a little large but without inducing obvious thermal effect.
@@
Simulations of JDE. A minimal model with $n$ JJs in parallel is used to explain the JDE 9, where only the first and second harmonic contributions in the CPR are considered. The supercurrent is approximatively described as $I(\varphi,\phi/\phi_{0})=\sum_{1}^{n}I_{n}sin(\varphi+2\pi\zeta_{n}\phi/\phi_{0})+\alpha_{n}I_{n}sin(2\varphi+4\pi\zeta_{n}\phi/\phi_{0})$, where $\zeta_{n}$ is the modified ratio and $\zeta_{1}$ = 0, $\beta_{2}$ = 1 ($\zeta_{1}$ = 0, $\zeta_{2}$ = 1, $\zeta_{3}$ = 0.23) for the simulation of our two (three) JJs in parallel. $\zeta_{3}$ represents the enclosed-area ratio of two corresponding SQUID patterns in device S2. Preliminary switching currents for different channels are estimated in both devices without considering the second harmonic contribution, in order to simulate CPR and JDE. For device S1, the parameters used for simulating $\Delta I_{c}$ are: $I_{1}$ = 75 nA, $I_{2}$ = 15 nA, $\alpha_{1}$ = 0.3 and $\alpha_{2}$ = 0.5, while for device S2, the parameters are: $I_{1}$ = 330 nA, $I_{2}$ = 120 nA, $I_{3}$ = 158 nA, $\alpha_{1}$ = 0.6, $\alpha_{2}$ = 0.1 and $\alpha_{3}$ = 0.2. The simulated $\phi/\phi_{0}$-dependent $\Delta I_{c}$ can approximatively explain the experimental data in both devices, as shown by red lines in Fig. 2d, i. The slight discrepancy between the simulation and experiment may arise from the omission of higher-order harmonics in the CPR, a reduction in critical current and a potential variation of $\alpha$ with increasing magnetic field, or other contributing factors.
@@
Secondly, there are several features supporting asymmetric edge supercurrents in our devices: (1) Without considering the higher harmonics of CPR, we calculated two edge supercurrent channels ($I_{1}$ and $I_{2}$) from the SQUID pattern in supplementary section III, and in device S1, the ratio between the supercurrents $I_{1}$:$I_{2}$ is approximately 3.5:1 ($I_{1}$:$I_{2}$:$I_{3}$ $\sim$ 6:1:2 for device S2); (2) After considering the second harmonic, from a simple simulation using a two-JJs model in supplementary Fig. S2e, the deviation of $I_{c}$ from zero at half flux quantum alone cannot support asymmetric interference channels. However, the coexistence of this deviation and the JDE can support asymmetric supercurrents, as shown in supplementary Fig. S2d, g, h, and device S1 satisfies this coexistence condition; (3) The position-dependent supercurrent density distribution shows a significantly asymmetric supercurrent distribution. However, it should be noted that such asymmetric supercurrent density distribution can also exist in a symmetric system 33; 34 due to the relatively imperfect experimental data.
@@
Discussion of switching power. From a practical perspective, the power consumption of devices should be carefully considered at low temperatures. Although the current cooling power of commercial dilution refrigerators generally reaches the microwatt level at 10 mK 66, a Joule excitation power higher than a picowatt (or nanowatt) may raise the temperature of the device or sensor above 10 mK (or 100 mK) when the mixing chamber is kept at 10 mK 67; 68. The self-heating effect is sometimes noticeable and can also be applied. For example, a microheater with a power output of a few picowatts can generate a temperature gradient as large as several millikelvin per micron, and can be further used to conduct Seebeck or Nernst effect measurements in 2D devices at millikelvin base temperatures 69. Therefore, ultra-low power consumption is essential not only for basic measurements but also for applications in large-scale superconducting quantum circuits, which need ultra-low temperatures to enhance the precision of quantum state measurements.
@@ Acknowledgments
We thank Jin-Guang Cheng for deep discussions. This work was supported by the National Natural Science Foundation of China (Grant Nos. 92065203, 12174430, U2032204, 11974395 and 12188101), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB33000000 and XDB33030000), the Beijing Nova Program (Grant No. Z211100002121144), the Synergetic Extreme Condition User Facility (SECUF), the Informatization Plan of Chinese Academy of Sciences (CAS-WX2021SF-0102), the Financial Support from Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302500), the China Postdoctoral Science Foundation (Grant No. 2021M703462 and 2021TQ0356) and the Center for Materials Genome. A portion of this work was carried out at the Synergetic Extreme Condition User Facility (SECUF).
@@
Competing interests: The authors declare no competing interests.

### v3

@@ Results
Furthermore, some features in Fig. 3 require further explanation. First, the disappearance of half-integer Shapiro steps at zero field (Fig. 3a, c) may be attributed to the relatively small contribution of the second harmonic in the CPR and the low microwave frequency 59. In device S1, a higher microwave frequency is utilized at zero field, and a small signal of half-integer steps appears in Supplementary Fig. 6d-f. No higher frequency is applied to detect half-integer steps at 0 T, considering the limitations of our microwave generator. Second, the integer Shapiro steps at 5.5 mT (3/2 $\phi_{0}$) appear weaker in Fig. 3b (or Supplementary Fig. 6c) compared to those at 0 T in Supplementary Fig. 6b. This is due to the application of a magnetic field at half the flux quantum, causing a reduction in the sharpness of the integer steps because of a decrease in the first harmonic of the CPR. Subsequently, the second harmonic or fractional Shapiro steps become visible at half the flux quantum (Fig. 3a), in agreement with theoretical predictions 9; 35 and experimental results 54; 24. Third, in device S2 (Fig. 3c), half-integer steps vanish at ±$\pm\phi_{0}/2$ due to the superposition of multiple SQUID patterns, which is slightly different from the single SQUID pattern observed in device S1. In the case of multiple SQUID patterns, the position of destructive interference for the first harmonic differs from that of a single SQUID pattern, leading to a notable contribution from the first harmonic and the absence of half-integer steps at ±$\pm\phi_{0}/2$ (see Supplementary Note 6).
Figure 4: Antisymmetric second harmonic transport of Ta2Pd3Te5 JJs. $\mathbf{a}$, $R_{w}$ versus $B_{z}$ at different currents. $\mathbf{b}$, $B_{z}$-dependent $R_{2w}$ at various currents. Typical pairs of antisymmetric peaks ($R_{2w}$) are marked by dashed lines in matching colors. $R_{2w}$ at 500 nA (black line) shows a similar behavior to $\Delta I_{c}$ (orange line). The green dashed line indicates the regions near the critical fields, while the other dashed line marks the areas near the oscillating peaks or valleys. $\mathbf{c}$, Typical $\left|R_{2w}\right|$ as a function of current at 20 mK for different oscillating peaks. The average $\left|R_{2w}\right|$ values are obtained using pairs of antisymmetric peaks near the dashed lines in $\mathbf{b}$. The error bars primarily originate from discrepancies between average $\left|R_{2w}\right|$ and the observed values. $\mathbf{d}$, Antisymmetric $R_{2w}$ and $\Delta I_{c}$ as a function of $B_{z}$ in device S1. The $R_{2w}$ is measured at $I_{ac}$ = 70 nA.
@@
The JDE is also studied as a function of microwave power. In Fig. 3f, switching current $I_{c+}^{0th}$ and $I_{c-}^{0th}$ of the 0th Shapiro step are extracted from the Fig. 3e. The quantity $\Delta I_{c}^{0th}=I_{c+}^{0th}-\left|I_{c-}^{0th}\right|$ (orange curve, see the detail in Supplementary Fig. 8) displays little variation and gradually decreases with increasing microwave power. This stable observation of the JDE under microwave irradiation represents one of the initial examples of such behavior and is a crucial feature in quantum information applications, where microwaves are commonly utilized 2.
Antisymmetric second harmonic transport
@@
As another type of non-reciprocal behavior, the second harmonic resistance ($R_{2w}$) is usually measured to characterize magnetochiral anisotropy (MCA) 60; 13; 25, whose strength is often defined by the coefficient $\gamma=\left|\frac{2R_{2w}}{R_{w}BI}\right|$. MCA typically exhibits antisymmetric behavior in $R_{2w}$ with respect to magnetic fields and is generally used to characterize noncentrosymmetric or chiral features. In normal conductors, MCA is usually a minor effect, generally due to the weak SOC. However, a significant enhancement of MCA is observed in noncentrosymmetric superconductors, potentially offering additional insights into factors such as disparities between the Fermi energy and superconducting gap, the Rashba parameter, and spin-orbit splitting, among other factors 60; 61. Moveover, the antisymmetric oscillations in $B$-dependent $R_{2w}$ have been observed in an ionic liquid-gated chiral nanotube exhibiting the Little-Parks effect 62, indicating its characteristic chiral symmetry.
Therefore, the second harmonic resistance in our supercurrent interferometer is measured to clarify its asymmetric features. In Fig. 4a, b, the first harmonic resistance ($R_{w}$) and $R_{2w}$ of device S2 are measured using lock-in techniques. Below $I_{ac}$ = 100 nA, a typical pair of antisymmetric peaks in $R_{2w}$ is observed at critical fields (near the green dashed line in Fig. 4b). This is consistent with the previously reported antisymmetric peaks of $R_{2w}$ with respect to $B$ in noncentrosymmetric superconductors 60; 13, suggesting the existence of asymmetry in our JJ. With increasing $I_{ac}$, the antisymmetric oscillations in $R_{2w}$ becomes more pronounced due to the quantum interference of asymmetric edge supercurrents, similar to the behavior observed in chiral nanotubes 62. At around 500 nA, the oscillating $R_{2w}$ (black line in Fig. 4b) is analogous to the $\Delta I_{c}$ (orange line), and the similar behavior of device S1 can be seen in Fig. 4d (see the detail in Supplementary Note 7). This similarity suggests the antisymmetric oscillations in $R_{2w}$ may be related to the polarity of JDE. At higher $I_{ac}$, the superconductivity is substantially suppressed and self-heating effects may begin to influence the second harmonic resistance, which is not further analyzed.
[…]
Discussion
@@
The Ta2Pd3Te5 edge interferometer reveals a significant JDE under small out-of-plane magnetic fields. The JDE is accompanied by the enhanced antisymmetric second harmonic transport, which deserves further research. Moreover, the presence of asymmetric Josephson critical current alongside fractional Shapiro steps demonstrates the importance of the higher harmonic content in the JDE. It is noted that power consumption requires attention if SDE/JDE is to be applied at low temperatures in the future, and further optimization is necessary not only regarding mechanism, but also in terms of device size. The Josephson diode efficiency can be further enhanced by concatenating interferometer loops. These findings offer a promising approach to exploring JDE with significant diode efficiency, ultra-low switching power at small magnetic fields and stable JDE under microwave irradiation, rendering them potentially valuable for practical applications.
Methods
@@
Device fabrication. Single crystals Ta2Pd3Te5 were prepared by the self-flux method 36. Ta2Pd3Te5 thin films were obtained through mechanical exfoliating bulk samples onto Si substrates with the 280-nm-thick SiO2 on top, and then coated with PMMA in a glove box at 4000 rpm for 60 s, followed by annealing at 100 C for 120 s. Multi-terminal electrical contacts were patterned after electron-beam exposure and subsequent development. Ti/Al (5 nm/60 nm) electrodes were deposited after Ar etching in order to remove the oxidized layer. After the lift-off step, the device was coated with hexagonal boron nitride (hBN) to prevent oxidation. The entire fabrication process took place in a nitrogen atmosphere glove box. The rectangular film, with a long side along the $b$-axis of the crystal, was easily obtained due to its quasi-1D nature 38. The dimensions of device S2 (device S1) are as follows: thickness 37 nm (16 nm), width 1.9 $\mu$m (1.3 $\mu$m), and length 245 nm (400 nm), respectively.
@@
Transport measurements under microwave radiation. The electrical transport measurements under microwave radiation were carried out using PSG-A Series Signal generator (Agilent E8254A). The main microwave frequency between 4.64 and 5.02 GHz was applied due to the large absorption effect of microwave for devices S1 and S2. The dissipation of microwave circuit existed during the measurement of device S2, while another microwave circuit for device S1 was good. So the applied power of microwave was a little large but without inducing obvious thermal effect.
@@
Simulations of JDE. A minimal model with $n$ JJs in parallel is used to explain the JDE 9, where only the first and second harmonic contributions in the CPR are considered. The supercurrent is approximatively described as $I(\varphi,\phi/\phi_{0})=\sum_{1}^{n}I_{n}sin(\varphi+2\pi\zeta_{n}\phi/\phi_{0})+\alpha_{n}I_{n}sin(2\varphi+4\pi\zeta_{n}\phi/\phi_{0})$, where $\zeta_{n}$ is the modified ratio and $\zeta_{1}$ = 0, $\zeta_{2}$ = 1 ($\zeta_{1}$ = 0, $\zeta_{2}$ = 1, $\zeta_{3}$ = 0.23) for the simulation of our two (three) JJs in parallel. $\zeta_{3}$ represents the enclosed-area ratio of two corresponding SQUID patterns in device S2. Preliminary switching currents for different channels are estimated in both devices without considering the second harmonic contribution, in order to simulate CPR and JDE. For device S1, the parameters used for simulating $\Delta I_{c}$ are: $I_{1}$ = 75 nA, $I_{2}$ = 15 nA, $\alpha_{1}$ = 0.3 and $\alpha_{2}$ = 0.5, while for device S2, the parameters are: $I_{1}$ = 330 nA, $I_{2}$ = 120 nA, $I_{3}$ = 158 nA, $\alpha_{1}$ = 0.6, $\alpha_{2}$ = 0.1 and $\alpha_{3}$ = 0.2. The simulated $\phi/\phi_{0}$-dependent $\Delta I_{c}$ can approximatively explain the experimental data in both devices, as shown by black lines in Fig. 2d, i. The slight discrepancy between the simulation and experiment may arise from the omission of higher-order harmonics in the CPR, a reduction in critical current and a potential variation of $\alpha$ with increasing magnetic field, or other contributing factors.
@@
Secondly, there are several features supporting asymmetric edge supercurrents in our devices: (1) Without considering the higher harmonics of CPR, we calculated two edge supercurrent channels ($I_{1}$ and $I_{2}$) from the SQUID pattern in Supplementary Note 3, and in device S1, the ratio between the supercurrents $I_{1}$:$I_{2}$ is approximately 3.5:1 ($I_{1}$:$I_{2}$:$I_{3}$ $\sim$ 6:1:2 for device S2); (2) After considering the second harmonic, from a simple simulation using a two-JJs model in Supplementary Fig. 2e, the deviation of $I_{c}$ from zero at half flux quantum alone cannot support asymmetric interference channels. However, the coexistence of this deviation and the JDE can support asymmetric supercurrents, as shown in Supplementary Fig. 2d, g, h, and device S1 satisfies this coexistence condition; (3) The position-dependent supercurrent density distribution shows a significantly asymmetric supercurrent distribution. However, it should be noted that such asymmetric supercurrent density distribution can also exist in a symmetric system 33; 34 due to the relatively imperfect experimental data.
@@
Discussion of switching power. From a practical perspective, the power consumption of devices should be carefully considered at low temperatures. Although the current cooling power of commercial dilution refrigerators generally reaches the microwatt level at 10 mK 67, a Joule excitation power higher than a picowatt (or nanowatt) may raise the temperature of the device or sensor above 10 mK (or 100 mK) when the mixing chamber is kept at 10 mK 68; 69. The self-heating effect is sometimes noticeable and can also be applied. For example, a microheater with a power output of a few picowatts can generate a temperature gradient as large as several millikelvin per micron, and can be further used to conduct Seebeck or Nernst effect measurements in 2D devices at millikelvin base temperatures 70. Therefore, ultra-low power consumption is essential not only for basic measurements but also for applications in large-scale superconducting quantum circuits, which need ultra-low temperatures to enhance the precision of quantum state measurements.
@@
Data availability
All relevant data are available from the authors. The data can also be found at the following link (http://dx.doi.org/10.6084/m9.figshare.26539759).
Code availability
DFT calculations can be reproduced using standard VASP packages. Simulations of JDE are fully described. The codes used in this study are available from the authors upon request.
@@ Acknowledgments
We thank Jin-Guang Cheng for deep discussions. This work was supported by the Beijing Natural Science Foundation (Grant No. JQ23022), the National Natural Science Foundation of China (Grant Nos. 92065203, 12174430, U2032204, 11974395, 12188101, U22A6005 and 12404154), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB33000000 and XDB33030000), the Beijing Nova Program (Grant No. Z211100002121144), the Synergetic Extreme Condition User Facility (SECUF), the Informatization Plan of Chinese Academy of Sciences (CAS-WX2021SF-0102), the Ministry of Science and Technology of China (Grants No. 2022YFA1403800), the Financial Support from Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302500), the China Postdoctoral Science Foundation (Grant No. 2021M703462 and 2021TQ0356) and the Center for Materials Genome. A portion of this work was carried out at the Synergetic Extreme Condition User Facility (SECUF).
@@
Competing interests The authors declare no competing interests.